OpenStudy (vortish):

solve and check x^2=3x+40

OpenStudy (anonymous):

x^2 - 3x - 40 = 0 (x-8)(x+5)=0 x=8 or x=-5 CHECK: 8^2 - 8(3) - 40 (-5)^2 - 3(-5) - 40

OpenStudy (anonymous):

@kwaldman answers are correct. ^